Two different companies, Vogel and Hatcher, entered into the following inventory transactions during December. Both companies use a perpetual inventory system.

  • December 3 – Vogel Corporation sold inventory on account to Hatcher Corp. for $240,000, terms 2/10, n/30. This inventory originally cost Vogel $160,000.
  • December 8 – Hatcher Corp. returned inventory to Vogel Corporation for a credit of $15,000. Vogel returned this inventory to inventory at its original cost of $10,000.
  • December 12 – Hatcher Corp. paid Vogel Corporation for the amount owed.

Required:

  1. Prepare the journal entries to record these transactions on the books of Vogel Corporation.
  2. What is the amount of net sales to be reported on Vogel Corporation’s income statement?
  3. What is the Vogel Corporation’s gross profit percentage?
Answer #1

Journal entries

Date account and explanation debit credit
Dec 3 Account receivable 240000
Sales revenue 240000
(To record sales)
Cost of goods sold 160000
Merchandise inventory 160000
(To record cost of goods sold)
Dec 8 Sales return and allowance 15000
Account receivable 15000
(To record sales return)
Merchandise inventory 10000
Cost of goods sold 10000
(To record cost of goods returned)
Dec 12 Cash (225000*98%) 220500
Sales discount 4500
Account receivable 250000
(To record cash received)

b) Net sales = 240000-15000-4500 = $220500

c) Gross profit = 220500-150000 = 70500

Gross profit percentage = 70500/220500 = 31.97%